OpenStudy (anonymous):

Find the equation of the line passing through the points (6, -1) and (5, 2). Write your answer in slope-intercept form.

OpenStudy (anonymous):

find the slope in your head. from 5 to 6 means right one unit. from 2 to -1 i down 3 units so slope is \[m=-3\]

OpenStudy (anonymous):

or you can use \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{-1-2}{6-5}=-3\]

OpenStudy (anonymous):

then use almighty point slope formula \[y-y_1=m(x-x_1)\] with \[m=-3,x_1=5,y_1=2\] to get \[y-2=-3(x-5)\] \[y-2=-3x+15\] \[y=-3x+17\]

OpenStudy (anonymous):

The answer of Satellite 73 is right

OpenStudy (anonymous):

Y=-3x + 17 is the answer of trykas question